OpenStudy (anonymous):

Simplify: 3x + 4y – 5(4x + 2) – 5y

OpenStudy (anonymous):

dang i really could swear i just did this exact problem

OpenStudy (anonymous):

lol can yu help me then

OpenStudy (anonymous):

\[ 3x+4y-20x-10-5y\] combine like terms to get

OpenStudy (anonymous):

-10 not positive 10

OpenStudy (anonymous):

we have \[3x-20x=-17x\] \[4y-5y=-y\] and of course 10 has to match to it is \[-17x-y-10\]

OpenStudy (anonymous):

i was to quick to draw myself

OpenStudy (anonymous):

*too

OpenStudy (anonymous):

3x+4y-20x-10-5y =3x-20x+4y-5y(grouping like terms) =-17x-y+10